A pump is required to lift 1000kg of water per minute from a well 12 m deep and eject it with a spee of 20m/s. what must be the power output of the pump?

Grade:

2 Answers

Arun
6 years ago
mass of the object=1000kg
time taken=1 min(60sec)
height attained =12 m
velocity attained= 20 m/s
∴ kinetic energy in the body 
E= 1/mv^2
E=1/2*1000*20^2
E=500*400
E=200000 J
∴power=energy/time
P=200000/60
P=3333.33 W
 
 
Regards
Arun (askIITians forum expert)
Priyansh Chittora
3 years ago
Mass - 1000kg
Height - 12 m
Velocity - 20 m/s
Work done - mgh = 1000 × 10 × 12 = 120000J
Work done per second - 120000/60 = 2000J
K.E gain by it per second - (1/2 mv^2)/60 = {1/2 × 1000 × (20)^2}/60 = 3333.33 J
Power output of the pump - 2000 + 3333.33 
= 5333.33 W or 5.33 KW